OpenStudy (anonymous):

Some investments in the stock market have earned 10% annually. At this rate, earnings can be found using the formula A=P(1.10)^n, where A is the total value of the investment, P is the initial value of the investment, and N is the number if years the money is invested. If $2,500 is invested in the stock market at this annual rate of return, what is the expected total value after 18 years? I feel like this should be easy, but i am lost.

OpenStudy (anonymous):

answer choices are: ~$49,500.00 ~$46,750.00 ~$13,899.79 ~$12,636.18

OpenStudy (anonymous):

Never mind guys i got it thanks though... its A

OpenStudy (anonymous):

This is just about understanding what information the text gives you: "$2,500 is invested in the stock market..." becomes your P, and "the expected total value" asks for A, after 18 years" is your n.

OpenStudy (anonymous):

The answer isn't option A...

OpenStudy (anonymous):

Plug in P and n into the equation given, and you'll get out A (in the equation) which is the answer.

OpenStudy (anonymous):

right... i took...oh... wait...

OpenStudy (anonymous):

i keep getting A...

OpenStudy (anonymous):

You have: A=P(1.10)^n P = 2500 n = 18 So: A = 2500(1.10)^18 = 2500*5.559... = 13899

OpenStudy (anonymous):

Did you do the same thing??

OpenStudy (anonymous):

um... yeah actualy

OpenStudy (anonymous):

Oh, maybe the order of your answers aren't actually A,B,C,D then.. c: